DuPont Analysis Last year, PJ's Ice Cream Parlors, Inc. reported an ROE = 12.6%. The firm's debt ratio was 44%, sales were $41 million, and the capital intensity ratio was .91 times. What is the net income for PJ's last year? (Do not round intermediate steps.)
Answer #1

capital intensity ratio = Total Assets / Sales

capital intensity ratio = 0.91

Sales = $41 m

=> Total Assets = capital intensity ratio * Sales = 0.91 * 41 = $37.31 m

Debt Ratio = Total Debt / Total Assets

Debt Ratio = 0.44

=> Total Debt = Debt Ratio * Total Assets = 0.44*37.31 = $16.42 m

Assets = Equity + Debt

=> Total Equity = Total Assets - Total Debt = 37.31 - 16.42 = $20.89 m

ROE = Net Income / Total Equity

ROE = 0.126

=> Net Income = ROE * Total Equity = 0.126*20.86 = $2.63 m